OpenStudy (anonymous):

can someone please help simplify sqr(288 x^7 y^3)

OpenStudy (anonymous):

Break it up first and it should be fairly straight forward:\[\sqrt{288}*\sqrt{x^7}*\sqrt{y^3}\]Hopefully you know some indicies laws and that the square root is the same as putting it to the power of a half, so you can write:\[12\sqrt{2}*x^{7/2}*y^{3/2}\]